# VMarker-Pro: Probabilistic 3D Human Mesh Estimation from Virtual Markers

### v3

@@ Abstract
Monocular 3D human mesh estimation faces challenges due to depth ambiguity and the complexity of mapping images to complex parameter spaces. Recent methods propose to use 3D poses as a proxy representation, which often lose crucial body shape information, leading to mediocre performance. Conversely, advanced motion capture systems, though accurate, are impractical for markerless wild images. Addressing these limitations, we introduce an innovative intermediate representation as virtual markers, which are learned from large-scale mocap data, mimicking the effects of physical markers. Building upon virtual markers, we propose VMarker, which detects virtual markers from wild images, and the intact mesh with realistic shapes can be obtained by simply interpolation from these markers. To address occlusions that obscure 3D virtual marker estimation, we further enhance our method with VMarker-Pro, a probabilistic framework that generates multiple plausible meshes aligned with images. This framework models the 3D virtual marker estimation as a conditional denoising process, enabling robust 3D mesh estimation. Our approaches surpass existing methods on three benchmark datasets, particularly demonstrating significant improvements on the SURREAL dataset, which features diverse body shapes. Additionally, VMarker-Pro excels in accurately modeling data distributions, significantly enhancing performance in occluded scenarios. Code and models are available at https://github.com/ShirleyMaxx/VMarker-Pro.
Index Terms:
Computer vision, 3D human mesh estimation, probabilistic modeling
Figure 1: Top: Mesh estimation results on examples with different body shapes. Pose2Mesh [1] which uses 3D skeletons as the intermediate representation fails to predict accurate shapes. Our virtual marker-based method VMarker obtains accurate estimates. Bottom: VMarker may fail when occlusion happens, e.g.  the occluded left arm is wrongly estimated. To handle the ambiguity, we propose a probabilistic framework VMarker-Pro which could generate multiple reasonable estimates aligning with the image.
@@ 1 Introduction
Monocular 3D human mesh estimation aims to estimate 3D positions of the mesh vertices that are on the body surface from an RGB image. The task has attracted a lot of attention from the computer vision and computer graphics communities [1, 2, 3, 4, 5, 6, 7, 8] because it can benefit many applications such as virtual reality [9]. Recently, deep learning-based methods [6, 1, 10] have significantly advanced the accuracy on benchmark datasets.
@@
Pioneer methods [11, 6] propose to regress the pose and shape parameters of the mesh models such as SMPL [12] directly from images, learning a one-to-one mapping from images to the parameter space. While straightforward, their performance usually suffers from several limitations. Firstly, the mapping from the image to the $SO(3)$ parameter space is highly non-linear and suffers from image-model misalignment [10]. Secondly, existing mesh datasets [13, 14, 15] are small and limited to simple laboratory environments due to the complex capturing process. The lack of sufficient training data severely limits its performance. Lastly, due to the inherent depth ambiguity in monocular estimation, multiple 3D meshes may correspond to the same 2D projection. As a result, these deterministic methods may produce incorrect estimates in the presence of occlusions.
@@
Recently, some works [16, 17] begin to formulate mesh estimation as a dense 3D keypoint detection task inspired by the success of volumetric pose estimation [18, 19, 20, 21, 22, 23]. For example, in [16, 17], the authors propose to regress the 3D positions of all vertices. However, it is computationally expensive because it has more than several thousand vertices. Moon et al.  [17] improves the efficiency by decomposing the 3D heatmaps into multiple 1D heatmaps at the cost of mediocre accuracy. Choi et al.  [1] propose to first detect a sparser set of skeleton joints in the images, from which dense 3D meshes are regressed by exploiting the mesh topology. The methods along this direction have attracted increasing attention [1, 10, 24] due to two reasons. First, the proxy task of 3D skeleton estimation can leverage the abundant 2D pose datasets which notably improves the accuracy. Second, mesh regression from the skeletons is efficient. However, important information about the body shapes is lost in extracting the 3D skeletons, which is largely overlooked. As a result, different types of body shapes, whether lean or obese, cannot be accurately estimated, e.g.  the results of Pose2Mesh [1] in Figure 1.
@@
The professional marker-based motion capture method MoSh [5] places physical markers on the body surface to extract meshes with accurate shapes under non-rigid motions. However, the physical markers limit the approach in laboratory environments. We are inspired to think whether we can identify a set of landmarks on the mesh as virtual markers, e.g.  , elbow and wrist, that can be detected from wild images, allowing us to recover accurate body shapes. The desired virtual markers should satisfy several requirements. First, the number of markers should be much smaller than that of the mesh vertices so that we can use volumetric representations to estimate their 3D positions efficiently. Second, the markers should capture the mesh topology so that the intact mesh can be accurately regressed from them. Third, the virtual markers have distinguishable visual patterns and can be detected from images.
@@
In this work, we present a learning algorithm based on archetypal analysis [25] to identify a subset of mesh vertices as the virtual markers that try to satisfy the above requirements to the best extent. Figure 2 shows that the learned virtual markers coarsely outline the body shape and pose which paves the way for estimating meshes with accurate shapes. Then we present a simple framework VMarker for 3D mesh estimation on top of the representation as shown in Figure 3 (top). It first detects the 3D positions of the virtual markers [18]. Then we recover the intact mesh simply by interpolating them. The interpolation weights are pre-trained in the representation learning step and will be adjusted by a light network based on the prediction confidences of the virtual markers for each image.
Although the virtual marker representation eases the mesh estimation task, the vulnerability of deterministic methods to depth ambiguity and occlusions underscores the need for probabilistic modeling. For example, in Figure 1 (bottom), VMarker wrongly estimates the occluded left arm. Expanding upon VMarker, we further develop VMarker-Pro, a probabilistic framework capable of estimating multiple plausible 3D meshes that are consistent with the given 2D observation, as depicted in Figure 3 (bottom). Motivated by recent breakthroughs in generative models, we utilize diffusion models [26] to enhance the diversity and robustness of our mesh predictions. Concretely, we formulate the estimation of 3D virtual markers as a conditional reverse denoising process. Based on the input 2D image and the predicted virtual markers from VMarker, we probabilistically model the 3D marker positions and finally obtain multiple reasonable 3D mesh estimates. For instance, VMarker-Pro could estimate a more naturally swinging left arm during walking than VMarker as shown in the bottom of Figure 1.
We extensively evaluate our approaches on three benchmark datasets. Both VMarker and VMarker-Pro consistently outperforms the state-of-the-art methods across the benchmarks. Notably, VMarker demonstrates remarkable gains on the SURREAL dataset [27], which has a variety of body shapes. Ablation studies further affirm the merits of virtual marker representation, particularly in estimating accurate shapes. Furthermore, VMarker-Pro stands out in its precise modeling of data distributions, markedly enhancing performance in occlusion-heavy scenarios. Finally, experiments show decent generalization ability of our methods which generate visually appealing results for wild images.
A preliminary version of this work was presented in CVPR 2023 [28]. Our new contributions are three-fold. First, we propose VMarker-Pro, which incorporates probabilistic modeling to handle the inherent ambiguity in monocular 3D mesh estimation. This yields a more flexible and robust human mesh distribution estimation. Second, VMarker-Pro outperforms state-of-the-arts (SOTA) on multiple benchmark datasets in terms of the accuracy of distribution. Finally, VMarker-Pro significantly improves the performance in scenarios with occlusions than VMarker and other SOTA methods, demonstrating its ability to resolve ambiguity.
@@ 2.1 Optimization-based mesh estimation
Before deep learning dominates this field, 3D human mesh estimation [5, 29, 15, 30, 31] is mainly optimization-based, which optimizes the parameters of the human mesh models to match the observations. For example, Loper et al.  [5] propose MoSh that optimizes the SMPL [12] parameters to align the mesh with the 3D marker positions. It is usually used to get GT 3D meshes for benchmark datasets because of its high accuracy. Later works propose to optimize the model parameters or mesh vertices based on 2D image cues [29, 15, 30, 31, 32]. They extract intermediate representations such as 2D skeletons from the images and optimize the mesh model by minimizing the discrepancy between the model projection and the intermediate representations such as the 2D skeletons. These methods are usually sensitive to initialization and suffer from local optimum.
Figure 2: Left: The learned virtual markers (blue balls) in the back and front views. The grey balls mean they are invisible in the front view. The virtual markers act similarly to physical body markers and approximately outline the body shape. Right: Mesh estimation results by our approach VMarker, from left to right are input image, estimated 3D mesh overlayed on the image, and three different viewpoints showing the estimated 3D mesh with our intermediate predicted virtual markers (blue balls), respectively.
@@ 2.2 Learning-based mesh estimation
Recently, most studies follow the learning-based framework and achieve promising results. Existing learning-based methods can roughly be divided into two categories: deterministic mesh estimation and probabilistic mesh estimation. A considerable amount of deterministic estimation work [11, 6, 33, 34, 35, 36, 17, 37, 38, 39, 40, 10, 1] aims at estimating a single 3D human mesh from a monocular image. However, considering the inherent depth ambiguity in monocular settings, an increasing number of works propose the probabilistic modeling framework which estimates multiple feasible solutions [41, 42, 43, 44, 45] to enhance the robustness of the model.
2.2.1 Deterministic mesh estimation
Pioneer works propose to use deep networks [11, 6, 33, 36] to regress the SMPL [12] parameters from image features. However, learning the mapping from the image space to the parameter space is highly non-linear [17]. In addition, they suffer from the misalignment between the meshes and image pixels [38]. These problems make it difficult to learn an accurate yet generalizable model.
@@ 2.2.1 Deterministic mesh estimation
2.2.2 Probabilistic mesh estimation
Despite the extensive exploration of multi-hypothesis 3D pose estimation [49, 50, 51, 52, 53, 54, 55], research on probabilistic 3D human mesh estimation remains relatively underdeveloped. Most recent approaches [41, 56, 42, 44, 45] predict distributions over SMPL [12] parameters conditioned on a 2D image. Biggs et al.  [41] expand upon HMR [6] to predict a categorical distribution over SMPL parameters and obtain multiple hypotheses discreetly. Sengupta et al.  [56] and ProPose [42] utilize the matrix Fisher distribution to model body pose rotation distributions. ProHMR [44] and HuManiFlow [43] employ Normalizing Flows [57, 58] to model the feasible 3D human model parameter distribution. More recently, HMDiff [45] adopts diffusion models [26] to generate multiple plausible human meshes.
Our work belongs to the learning-based class, offering both deterministic and probabilistic approaches. The deterministic approach, VMarker, introduces virtual markers as a novel intermediate representation, improving upon traditional physical markers or skeletons to better disambiguate pose and shape with more expressiveness. Our probabilistic framework, VMarker-Pro, employs diffusion models to produce diverse and plausible virtual marker estimates, from which multiple full mesh estimates that align with the image cues can be obtained.
2.3 Diffusion models
Denoising Diffusion Probabilistic Models (DDPMs) [59, 26] are a class of generative models that garner attention for their ability to generate high-quality images [26]. DDPMs employ a forward process where noise is incrementally added to the data until a noise distribution is reached, paired with a reverse process that incrementally denoises to generate data from noise. The strength of DDPMs lies in their iterative refinement, effectively capturing the target data distribution and enabling the synthesis of new samples. With the development of acceleration [26, 60] and enhancement [61, 62], the versatility of DDPMs has been demonstrated across various tasks beyond image generation [26, 60], contributing to innovations in video generation [63, 64, 65], semantic segmentation [66], motion generation [67, 68, 69], etc.
In the realm of 3D mesh estimation, existing methods [45] apply DDPMs to model a set of downsampled mesh vertices, which is hard to detect from images [17, 28]. Our approach utilizes DDPMs to model the distribution of virtual markers, an efficient and expressive intermediate representation. This reduces the complexity of the modeling, resulting in more robust and accurate estimations.
@@ 3 Method
In this section, we describe the details of our approach. First, Section 3.1 introduces how we learn the virtual marker representation from mocap data. Based on the virtual marker, we present the deterministic framework VMarker for mesh estimation from an image and its training details in Section 3.2. Then we extend VMarker to a probabilistic mesh estimation framework, VMarker-Pro in Section 3.3 and introduce its training details as well as the inference strategy.
@@ 3.1 The virtual marker representation
We represent a 3D mesh by a vector of vertex positions $\mathbf{x}\in\mathbb{R}^{3M}$ where $M$ is the number of mesh vertices. Denote a mocap dataset such as [13] with $N$ meshes as $\overset{\frown}{\mathbf{X}}=[\mathbf{x}_{1},\,...,\,\mathbf{x}_{N}]\in\mathbb{R}^{3M\times N}$. To unveil the latent structure among vertices, we reshape it to $\mathbf{X}\in\mathbb{R}^{3N\times M}$ with each column $\mathbf{x}_{i}\in\mathbb{R}^{3N}$ representing all possible positions of the $i^{\text{th}}$ vertex in the dataset [13].
The mesh $\mathbf{X}$ can be approximated by a low-rank matrix because the mesh representation is smooth and redundant where some vertices can be accurately reconstructed by others. Inspired by this, we aim to learn $K$ virtual markers from $M$ mesh vertices $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{K}]\in\mathbb{R}^{3N\times K}$, as shown in Figure 2, that try to satisfy the following two requirements to the greatest extent. First, they can accurately reconstruct the intact mesh $\mathbf{X}$ by their linear combinations: $\mathbf{X}=\mathbf{Z}\mathbf{A}$, where $\mathbf{A}\in\mathbb{R}^{K\times M}$ is a coefficient matrix that encodes the spatial relationship between the virtual markers and the mesh vertices. Second, they should have distinguishable visual patterns in images so that they can be easily detected from images. Ideally, they can be on the body surface as the meshes. While it seems natural to apply singular value decomposition (SVD) [70] to $\mathbf{X}$ to compute the eigenvectors as virtual markers for reconstructing others, there is no guarantee that the virtual markers correspond to the mesh vertices, making them difficult to be detected from images.
To that end, we apply archetypal analysis [25, 71] to learn $\mathbf{Z}$ by minimizing a reconstruction error with two additional constraints: (1) each vertex $\mathbf{x}_{i}$ can be reconstructed by convex combinations of $\mathbf{Z}$, and (2) each marker $\mathbf{z}_{i}$ should be convex combinations of the mesh vertices $\mathbf{X}$:
@@
where $\mathbf{A}=[\bm{\alpha}_{1},...,\bm{\alpha}_{M}]\in\mathbb{R}^{K\times M}$, each $\bm{\alpha}$ resides in the simplex ${\Delta}_{K}\triangleq\{\bm{\alpha}\in\mathbb{R}^{K}\,\mathrm{s.t.}\,\bm{\alpha}\succeq 0\,\text{and}\,{||\bm{\alpha}||}_{1}=1\}$, and $\mathbf{B}=[\bm{\beta}_{1},...,\bm{\beta}_{K}]\in\mathbb{R}^{M\times K}$, $\bm{\beta}_{j}\in{\Delta}_{M}$. We adopt Active-set algorithm [71] to solve objective (Eq. 1) and obtain the learned virtual markers $\mathbf{Z}=\mathbf{X}\mathbf{B}\in\mathbb{R}^{3N\times K}$. As shown in [25, 71], the two constraints encourage the virtual markers $\mathbf{Z}$ to unveil the latent structure among vertices, therefore they learn to be close to the extreme points of the mesh and located on the body surface as much as possible.
Figure 3: Overview of our deterministic (top) and probabilistic (bottom) framework. Top: The deterministic framework VMarker. Given an input image, it first estimates the 3D virtual marker (VM) positions $\mathbf{P}$. Then we update the coefficient matrix $\mathbf{A}$ based on the estimation confidence scores $\mathbf{C}$ of the VM. Finally, the complete human mesh can be simply recovered by linear multiplication $\mathbf{M}=\mathbf{P}\mathbf{A}$. Bottom: The probabilistic framework VMarker-Pro. Building upon VMarker  we formulate the 3D VM estimation as a reverse denoising process using diffusion models. We introduce a denoiser that learns to remove the added noise from a noisy 3D VM $\hat{\mathbf{P}}(t)$ during training, conditioned on the 2D image feature $\mathbf{F}$ and the estimated 2D VMs $\mathbf{P}^{2d}$, i.e.  the first two dimensions of $\mathbf{P}$. At inference, we iteratively removes noise from random noise $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ according to Eq. 16, yielding a denoised 3D VMs $\mathbf{P}(0)$. Finally, we obtain the estimated mesh $\mathbf{M}=\mathbf{P}(0)\mathbf{A}$. VMarker-Pro can generate multiple feasible solutions aligned with 2D image cues during inference.
Post-processing. Since human body is left-right symmetric, we adjust $\mathbf{Z}$ to reflect the property. We first replace each $\mathbf{z}_{i}\in\mathbf{Z}$ by its nearest vertex on the mesh and obtain $\widetilde{\mathbf{Z}}\in\mathbb{R}^{3\times K}$. This is done on a template mesh which is in a canonical pose. This step allows us to compute the left or right counterpart of each marker. Then we replace the markers in the right body with the symmetric vertices in the left body and obtain the symmetric markers $\widetilde{\mathbf{Z}}^{sym}\in\mathbb{R}^{3\times K}$. Finally, we update $\mathbf{B}$ and $\mathbf{A}$ by minimizing $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ subject to $\widetilde{\mathbf{Z}}^{sym}=\mathbf{X}\widetilde{\mathbf{B}}^{sym}$ over all $N$ mesh data. More details are elaborated in the supplementary.
Figure 2 shows the virtual markers learned on the mocap dataset [13] after post-processing. They are similar to the physical markers and approximately outline the body shape which agrees with our expectations. They are roughly evenly distributed on the surface of the body, and some of them are located close to the body keypoints, which have distinguishable visual patterns to be accurately detected. Table I shows the reconstruction errors of using original markers $\mathbf{X}\mathbf{B}$ and the symmetric markers $\mathbf{X}\widetilde{\mathbf{B}}^{sym}$. Both can reconstruct meshes accurately.
TABLE I: The reconstruction errors using the original and the symmetric sets of markers on the H3.6M dataset [13], respectively. The errors are small indicating that they are sufficiently expressive and can reconstruct all vertices accurately.
@@
3.2 Deterministic mesh estimation framework
3.2.1 VMarker
On top of the virtual markers, we present a simple yet effective framework VMarker for end-to-end 3D human mesh estimation from a single image. As shown in Figure 3 (top), it consists of two branches (yellow and orange blocks). The first branch (yellow block) uses a volumetric CNN [18] to estimate the 3D virtual marker (VM) positions $\mathbf{P}$, and the second branch (orange block) reconstructs the full mesh $\mathbf{M}$ by predicting a coefficient matrix $\mathbf{A}$:
$\mathbf{M}=\mathbf{P}\mathbf{A}.$ (2)

### v4

# 3D Human Mesh Estimation from Virtual Markers

@@ Abstract
Inspired by the success of volumetric 3D pose estimation, some recent human mesh estimators propose to estimate 3D skeletons as intermediate representations, from which, the dense 3D meshes are regressed by exploiting the mesh topology. However, body shape information is lost in extracting skeletons, leading to mediocre performance. The advanced motion capture systems solve the problem by placing dense physical markers on the body surface, which allows to extract realistic meshes from their non-rigid motions. However, they cannot be applied to wild images without markers. In this work, we present an intermediate representation, named virtual markers, which learns 64 landmark keypoints on the body surface based on the large-scale mocap data in a generative style, mimicking the effects of physical markers. The virtual markers can be accurately detected from wild images and can reconstruct the intact meshes with realistic shapes by simple interpolation. Our approach outperforms the state-of-the-art methods on three datasets. In particular, it surpasses the existing methods by a notable margin on the SURREAL dataset, which has diverse body shapes. Code is available at https://github.com/ShirleyMaxx/VirtualMarker.
@@ 1 Introduction
Figure 1: Mesh estimation results on four examples with different body shapes. Pose2Mesh [7] which uses 3D skeletons as the intermediate representation fails to predict accurate shapes. Our virtual marker-based method obtains accurate estimates.
@@
3D human mesh estimation aims to estimate the 3D positions of the mesh vertices that are on the body surface. The task has attracted a lot of attention from the computer vision and computer graphics communities [3, 43, 30, 35, 51, 19, 25, 37, 27, 10] because it can benefit many applications such as virtual reality [15]. Recently, the deep learning-based methods [19, 7, 29] have significantly advanced the accuracy on the benchmark datasets.
@@
The pioneer methods [51, 19] propose to regress the pose and shape parameters of the mesh models such as SMPL [36] directly from images. While straightforward, their accuracy is usually lower than the state-of-the-arts. The first reason is that the mapping from the image features to the model parameters is highly non-linear and suffers from image-model misalignment [29]. Besides, existing mesh datasets [16, 54, 38, 28] are small and limited to simple laboratory environments due to the complex capturing process. The lack of sufficient training data severely limits its performance.
@@
Recently, some works [26, 39] begin to formulate mesh estimation as a dense 3D keypoint detection task inspired by the success of volumetric pose estimation [47, 50, 65, 44, 59, 45]. For example, in [26, 39], the authors propose to regress the 3D positions of all vertices. However, it is computationally expensive because it has more than several thousand vertices. Moon and Lee [39] improve the efficiency by decomposing the 3D heatmaps into multiple 1D heatmaps at the cost of mediocre accuracy. Choi et al. [7] propose to first detect a sparser set of skeleton joints in the images, from which the dense 3D meshes are regressed by exploiting the mesh topology. The methods along this direction have attracted increasing attention [7, 29, 55] due to two reasons. First, the proxy task of 3D skeleton estimation can leverage the abundant 2D pose datasets which notably improves the accuracy. Second, mesh regression from the skeletons is efficient. However, important information about the body shapes is lost in extracting the 3D skeletons, which is largely overlooked previously. As a result, different types of body shapes, such as lean or obese, cannot be accurately estimated (see Figure 1).
@@
The professional marker-based motion capture (mocap) method MoSh [35] places physical markers on the body surface and explore their subtle non-rigid motions to extract meshes with accurate shapes. However, the physical markers limit the approach to be used in laboratory environments. We are inspired to think whether we can identify a set of landmarks on the mesh as virtual markers, e.g., elbow and wrist, that can be detected from wild images, and allow to recover accurate body shapes? The desired virtual markers should satisfy several requirements. First, the number of markers should be much smaller than that of the mesh vertices so that we can use volumetric representations to efficiently estimate their 3D positions. Second, the markers should capture the mesh topology so that the intact mesh can be accurately regressed from them. Third, the virtual markers have distinguishable visual patterns so that they can be detected from images.
@@
In this work, we present a learning algorithm based on archetypal analysis [12] to identify a subset of mesh vertices as the virtual markers that try to satisfy the above requirements to the best extent. Figure 2 shows that the learned virtual markers coarsely outline the body shape and pose which paves the way for estimating meshes with accurate shapes. Then we present a simple framework for 3D mesh estimation on top of the representation as shown in Figure 3. It first learns a 3D keypoint estimation network based on [47] to detect the 3D positions of the virtual markers. Then we recover the intact mesh simply by interpolating them. The interpolation weights are pre-trained in the representation learning step and will be adjusted by a light network based on the prediction confidences of the virtual markers for each image.
We extensively evaluate our approach on three benchmark datasets. It consistently outperforms the state-of-the-art methods on all of them. In particular, it achieves a significant gain on the SURREAL dataset [53] which has a variety of body shapes. Our ablation study also validates the advantages of the virtual marker representation in terms of recovering accurate shapes. Finally, the method shows decent generalization ability and generates visually appealing results for the wild images.
@@ 2.1 Optimization-based mesh estimation
Before deep learning dominates this field, 3D human mesh estimation [35, 2, 28, 41, 60] is mainly optimization-based, which optimizes the parameters of the human mesh models to match the observations. For example, Loper et al. [35] propose MoSh that optimizes the SMPL parameters to align the mesh with the 3D marker positions. It is usually used to get GT 3D meshes for benchmark datasets because of its high accuracy. Later works propose to optimize the model parameters or mesh vertices based on 2D image cues [2, 28, 41, 60, 11]. They extract intermediate representations such as 2D skeletons from the images and optimize the mesh model by minimizing the discrepancy between the model projection and the intermediate representations such as the 2D skeletons. These methods are usually sensitive to initialization and suffer from local optimum.
Figure 2: Left: The learned virtual markers (blue balls) in the back and front views. The grey balls mean they are invisible in the front view. The virtual markers act similarly to physical body markers and approximately outline the body shape. Right: Mesh estimation results by our approach, from left to right are input image, estimated 3D mesh overlayed on the image, and three different viewpoints showing the estimated 3D mesh with our intermediate predicted virtual markers (blue balls), respectively.
@@ 2.2 Learning-based mesh estimation
Recently, most works follow the learning-based framework and have achieved promising results. Deep networks [51, 19, 25, 37, 27] are used to regress the SMPL parameters from image features. However, learning the mapping from the image space to the parameter space is highly non-linear [39]. In addition, they suffer from the misalignment between the meshes and image pixels [62]. These problems make it difficult to learn an accurate yet generalizable model.
@@
Our work belongs to the learning-based class and is related to works that use physical markers or skeletons as intermediate representations. But different from them, we propose a novel intermediate representation, named virtual markers, which is more expressive to reduce the ambiguity in pose and shape estimation than body skeletons and can be applied to wild images.
@@ 3 Method
In this section, we describe the details of our approach. First, Section 3.1 introduces how we learn the virtual marker representation from mocap data. Then we present the overall framework for mesh estimation from an image in Section 3.2. At last, Section 3.3 discusses the loss functions and training details.
@@ 3.1 The virtual marker representation
We represent a mesh by a vector of vertex positions $\mathbf{x}\in\mathbb{R}^{3M}$ where $M$ is the number of mesh vertices. Denote a mocap dataset such as [16] with $N$ meshes as $\overset{\frown}{\mathbf{X}}=[\mathbf{x}_{1},\,...,\,\mathbf{x}_{N}]\in\mathbb{R}^{3M\times N}$. To unveil the latent structure among vertices, we reshape it to $\mathbf{X}\in\mathbb{R}^{3N\times M}$ with each column $\mathbf{x}_{i}\in\mathbb{R}^{3N}$ representing all possible positions of the $i^{\text{th}}$ vertex in the dataset [16].
The rank of $\mathbf{X}$ is smaller than $M$ because the mesh representation is smooth and redundant where some vertices can be accurately reconstructed by the others. While it seems natural to apply PCA [18] to $\mathbf{X}$ to compute the eigenvectors as virtual markers for reconstructing others, there is no guarantee that the virtual markers correspond to the mesh vertices, making them difficult to be detected from images. Instead, we aim to learn $K$ virtual markers $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{K}]\in\mathbb{R}^{3N\times K}$ that try to satisfy the following two requirements to the greatest extent. First, they can accurately reconstruct the intact mesh $\mathbf{X}$ by their linear combinations: $\mathbf{X}=\mathbf{Z}\mathbf{A}$, where $\mathbf{A}\in\mathbb{R}^{K\times M}$ is a coefficient matrix that encodes the spatial relationship between the virtual markers and the mesh vertices. Second, they should have distinguishable visual patterns in images so that they can be easily detected from images. Ideally, they can be on the body surface as the meshes.
We apply archetypal analysis [12, 4] to learn $\mathbf{Z}$ by minimizing a reconstruction error with two additional constraints: (1) each vertex $\mathbf{x}_{i}$ can be reconstructed by convex combinations of $\mathbf{Z}$, and (2) each marker $\mathbf{z}_{i}$ should be convex combinations of the mesh vertices $\mathbf{X}$:
@@
where $\mathbf{A}=[\bm{\alpha}_{1},...,\bm{\alpha}_{M}]\in\mathbb{R}^{K\times M}$, each $\bm{\alpha}$ resides in the simplex ${\Delta}_{K}\triangleq\{\bm{\alpha}\in\mathbb{R}^{K}\,\mathrm{s.t.}\,\bm{\alpha}\succeq 0\,\text{and}\,{||\bm{\alpha}||}_{1}=1\}$, and $\mathbf{B}=[\bm{\beta}_{1},...,\bm{\beta}_{K}]\in\mathbb{R}^{M\times K}$, $\bm{\beta}_{j}\in{\Delta}_{M}$. We adopt Active-set algorithm [4] to solve objective (1) and obtain the learned virtual markers $\mathbf{Z}=\mathbf{X}\mathbf{B}\in\mathbb{R}^{3N\times K}$. As shown in [12, 4], the two constraints encourage the virtual markers $\mathbf{Z}$ to unveil the latent structure among vertices, therefore they learn to be close to the extreme points of the mesh and located on the body surface as much as possible.
Figure 3: Overview of our framework. Given an input image $\mathbf{I}$, it first estimates the 3D positions $\hat{\mathbf{P}}$ of the virtual markers. Then we update the coefficient matrix $\hat{\mathbf{A}}$ based on the estimation confidence scores $\mathbf{C}$ of the virtual markers. Finally, the complete human mesh can be simply recovered by linear multiplication $\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}$.
@@
Table 1: The reconstruction errors using the original and the symmetric sets of markers on the H3.6M dataset [16], respectively. The errors are small indicating that they are sufficiently expressive and can reconstruct all vertices accurately.
Post-processing. Since human body is left-right symmetric, we adjust $\mathbf{Z}$ to reflect the property. We first replace each $\mathbf{z}_{i}\in\mathbf{Z}$ by its nearest vertex on the mesh and obtain $\widetilde{\mathbf{Z}}\in\mathbb{R}^{3\times K}$. This step allows us to compute the left or right counterpart of each marker. Then we replace the markers in the right body with the symmetric vertices in the left body and obtain the symmetric markers $\widetilde{\mathbf{Z}}^{sym}\in\mathbb{R}^{3\times K}$. Finally we update $\mathbf{B}$ and $\mathbf{A}$ by minimizing $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ subject to $\widetilde{\mathbf{Z}}^{sym}=\mathbf{X}\widetilde{\mathbf{B}}^{sym}$. More details are elaborated in the supplementary.
Figure 2 shows the virtual markers learned on the mocap dataset [16] after post-processing. They are similar to the physical markers and approximately outline the body shape which agrees with our expectations. They are roughly evenly distributed on the surface of the body, and some of them are located close to the body keypoints, which have distinguishable visual patterns to be accurately detected. Table 1 shows the reconstruction errors of using original markers $\mathbf{X}\mathbf{B}$ and the symmetric markers $\mathbf{X}\widetilde{\mathbf{B}}^{sym}$. Both can reconstruct meshes accurately.
3.2 Mesh estimation framework
On top of the virtual markers, we present a simple yet effective framework for end-to-end 3D human mesh estimation from a single image. As shown in Figure 3, it consists of two branches. The first branch uses a volumetric CNN [47] to estimate the 3D positions $\hat{\mathbf{P}}$ of the markers, and the second branch reconstructs the full mesh $\hat{\mathbf{M}}$ by predicting a coefficient matrix $\hat{\mathbf{A}}$:
$\small\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}.$ (2)